OpenStudy (anonymous):

Solve y equals two times x plus three for x. x equals y divided by negative two minus three x equals y plus three divided by two x equals y minus three divided by negative two x equals y divided by three plus two

OpenStudy (austinl):

\[y = 2x+3\] Subtract 3 from both sides. \[y-3 = (2x+3)-3 = 2x\] \[y-3=2x\] Divide both sides by 2. \[\frac{y-3}{2} = x\]

OpenStudy (austinl):

Gee thanks for the medal or even a "thanks"

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

@austinL